Elasticity of Demand: A public health issue that impacts the cost of health care individually and on a societal level is that of obesity. One of the public policy tools that policy makers have is to impose taxes to impact prices and ultimately change consumption levels of products that contribute to obesity. Sugar soda (or pop, as many say in Minnesota) is one of these products. If policy makers know the elasticity of demand for a product, they may be able to choose a tax level to raise the price enough to impact the choice to purchase soda. During a three-month period, scanner data (from grocery stores, for example) was collected on soda purchases. The price of a case of soda (24 cans) increased from $5.99 to $7.99. As a consequence, during this three-month period the total quantity of soda purchased in the region decreased from 4,395 cases per day to 2,488 cases per day. 


1
Expert's answer
2021-11-07T19:44:40-0500

a)

"price \space elasticity=\frac{\frac{Q_2-Q_1}{\frac{Q_2+Q_1}{2}}}
{\frac{2747-4395}{\frac{2747+4395}{2}}}\\
=\frac{\frac{7.35-5.99}{\frac{7.352+5.99}{2}}}
{\frac{P_2-P_1}{\frac{P_2+P_1}{2}}}\\"

"=--2.2637"

(b) The price elasticity of demand is defined as the change in quantity demanded as a proportion of a price increase. In this situation, we can state that a unit percent rise in the price of a soda case resulted in a 2.2634 percent drop in the quantity demanded of soda cases.

Demand is indeed responsive at the given pricing and volumes. If the elasticity is more than one, the demand is said to be elastic. This suggests that demand falls faster than the price rises, indicating that demand is responsive to price changes. If the elasticity is less than 2, the demand is said to be inelastic, meaning that demand falls faster than price rises (in which case, the demand is non-responsive). Also, if the elasticity is equal to 1, we can say that the demand is unitary elastic, meaning that it falls in lockstep with the price increase.

(c) Because demand is elastic/responsive, applying a soda tax will result in the desired behavior change. The tax's goal is to dramatically limit the amount of soda drunk, as soda use is linked to health issues.

In this instance, even a tiny tax would result in a lower quantity demanded than the price rise (or increase in price due to the tax). As a result, applying a tax would have a greater impact because demand is elastic. This means that the change in quantity demanded in percentage terms would be greater than the change in price owing to the tax in percentage terms for any given tax.